# Efficient Multi Subject Visual Reconstruction from fMRI Using Aligned Representations

### v1

@@ A.2 Projection to 4096 Dimensions with Stable Diffusion XL
A.3 Experiments with different architectures
We also present results (Table 5) where we compare normal training vs AAMax training on the MindEye2 architecture. We can observe that performance does not decline at all as we move to different architectures, as AAMax is agnostic of the size, dimensionality and the components of fMRI reconstruction pipelines.
Table 5: Performance metrics when fine-tuning for Subject 2 with MindEye2’s architecture. Regular FT stands for regular fine-tuning (row 2). AA stands for adapter alignment (row 3). These results are computed when training on the entire dataset for 200 epochs.
Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
Subj1 Pretraining 0.291 0.398 95.81% 98.55% 94.21% 90.15% 0.679 0.388
Subj2 Normal FT 0.108 0.347 93.25% 97.61% 90.57% 87.04% 0.844 0.543
Subj2 AAMax FT 0.218 0.352 95.37% 97.78% 91.97% 88.76% 0.711 0.531
Appendix B Common Space Empirical Analysis
@@
B.1 Eigenvector Analysis
To first investigate the structural alignment of embeddings across subjects, we analyzed the subject-wise cosine similarity of the principal eigenvectors in the embedding spaces produced by two techniques: the standard end-to-end training (Normal) and our proposed approach (AAMax). For both techniques, we extracted the top five principal eigenvectors from the test data for each subject and computed the cosine similarity between the eigenvectors from different subjects.
@@ B.1 Eigenvector Analysis
B.2 Mean Squared Error
We compare the end-to-end training approach with AAMax by evaluating the MSE between subjects using shared images after training. In the end-to-end model, we observe relatively high MSE between subjects, indicating poor alignment in the common space. In contrast, the MSE for AAMax is an order of magnitude lower, demonstrating a much stronger alignment of subject-specific embeddings. This result suggests that AAMax significantly enhances alignment across subjects, making it far more effective than the normal approach in bringing embeddings closer together in the common space.
Figure 10: Comparing Mean Squared Error for the shared images across subjects in both training techniques. The MSE is compared after training is completed in both techniques. MSE for AAMax is an order of magnitude lower than normal end-to-end training
B.3 k-Nearest Neighbor Consistency
We also evaluate the k-nearest neighbors (k-NN) consistency between subjects for shared images after training, using 50 nearest neighbors as the metric. In the end-to-end model, the k-NN consistency is relatively low, indicating that the nearest neighbors of embeddings from one subject do not align well with those from another subject. However, with AAMax, the k-NN consistency significantly improves, demonstrating much better alignment of subject-specific embeddings. AAMax consistently outperforms the normal approach, highlighting its effectiveness in preserving semantic relationships across subjects in the common space.
Figure 11: Comparing kNN consistency @ 50 for the shared images across subjects in both training techniques. The kNN-C is compared after training is completed in both techniques. AAMax significantly outperforms end-to-end training.
Appendix C Emergence of Common Concepts: Additional Details
The existence of a semantic common representation space is suggested from the emergence of semantic categories in this space. For this, we extracted the most significant left singular vectors from the ridge regression weight matrix. This matrix was extracted from the pre-trained MindEye2 (Scotti et al., 2024) model which was trained using the 1024-dimensional common space. Subsequently, we projected the 1024-dimensional embeddings of each image onto these vectors and sampled from the two extremes (set 1 and set 3) and the middle (set 2). In order to obtain an objective classification, we presented these sets of images to GPT-4o (Achiam et al., 2023) and asked it to find the semantic categories that are well represented in set 1, somewhat represented in set 2 and not represented in set 3. Results for the first five singular vectors are shown in Table 6. As is evident, images in the principal dimensions of the representation space have a common theme.
@@ Appendix C Emergence of Common Concepts: Additional Details
Previous studies, such as those by (Huth et al., 2012), have indicated that animal categories, including humans, are represented distinctly from non-animal categories within the brain. In our study, we observed that the common representational space also preserves these semantic properties. As demonstrated in Figure 12, after projecting the fMRI signals of Subject 2 into the common space and onto the first two singular vectors, we can see the animal vs. non-animal distinction, consistent with previous hypotheses. Extending this analysis across additional dimensions, we observed similar separations between other hypothesized semantic categories of ”Social interaction,” ”man-made vs nature,” and ”biological vs non-biological.”
@@
To determine if dominant semantic categories were present within these sets, we employed ChatGPT with the query: “Is there a dominant semantic category to which the majority of images in a set belong?” The methodology for defining semantic categories was consistent with that used in (Huth et al., 2012). Table 6 illustrates our findings.
After projecting the fMRI signals of individual images from Subject 2 into the common space and onto the first singular vectors, we observe that certain concepts are strongly represented along these dimensions defined by the singular vectors, aligning with previous hypotheses (Figures 13, 14).
@@
Appendix D NP-Hardness Proof
@@ D.1 Problem Definition
Definition D.1 (Bin-Mapping Problem).
@@
D.2 Reduction from Set Cover
@@
Definition D.2 (Set Cover Problem).
@@
Appendix E Proof of Approximation Ratio
In this section, we prove that the gap function $\text{Gap}(S)$, minimized by our greedy heuristic, satisfies the submodularity property. Submodularity guarantees the diminishing returns property, which forms the basis for the $1-\frac{1}{e}$ approximation ratio achieved by the greedy algorithm for submodular optimization  (Fisher et al., 1978).
E.1 Gap Function Definition
@@
E.2 Submodularity Property
A function $f(S)$ is submodular if it satisfies the diminishing returns property  (Schrijver et al., 2003):
@@ E.4 Monotonicity of $\text{Gap}(S)$
E.5 Conclusion
Since $\text{Gap}(S)$ is both submodular and monotone, the greedy algorithm applied to minimize the gap function achieves a $1-\frac{1}{e}$ approximation ratio  (Fisher et al., 1978).
Appendix F Greedy Image Selection Algorithm
@@
Appendix G Training Time
Table 7: Comparison of Training Times for Normal and AAMax Training Methods. Numbers are averaged over all 3 fine tune subjects (2,5,7). The Adapters are first allowed to overfit by training them upto 400 epochs using MSE loss. This takes around a minute. Then end-to-end alignment is performed. It is important to note that besides the addition of the MSE Loss at the adapter level, both pipelines are running the same architecture and data and performance is thus very similar. The final time given is the time to train up to 160 epochs. But AAMax training can outperform 160 epochs of Normal training with just one epoch of end-to-end alignment.
@@ Appendix G Training Time
Normal Training - 221.4 seconds 590.4 minutes 590.4 minutes
@@
Appendix H Additional Experimental Results
H.1 Best Candidate Image Selection: Additional Evaluation
We conducted several experiments to demonstrate that our image selection strategy outperforms the baseline method in a limited data setting. Below we present our experimental results.
Table 8: Comparison of performance on our data selection strategy with varying number of dimensions (singular values) for Subject 2. Our method has not been optimized for low-level metrics, which are shown in gray for reference. 150 images were used for all the experiments in this table.
Low-Level High-Level
Dimensions PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
10 0.099 0.254 74.27% 83.55% 75.20% 81.03% 0.835 0.520
20 0.105 0.220 73.06% 83.33% 77.23% 79.98% 0.828 0.508
30 0.119 0.288 74.51% 82.11% 75.41% 79.12% 0.836 0.516
40 0.103 0.212 72.22% 83.35% 77.04% 79.44% 0.839 0.519
50 0.107 0.217 73.73% 82.76% 76.63% 78.75% 0.845 0.524
60 0.101 0.202 72.91% 83.49% 75.97% 79.97% 0.829 0.519
Table 9: Comparison of performance on our data selection strategy with varying number of dimensions (singular values) for Subject 5. Our method has not been optimized for low-level metrics, which are shown in gray for reference. 150 images were used for all the experiments in this table.
Low-Level High-Level
Dimensions PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
10 0.138 0.292 75.20% 83.81% 78.62% 83.23% 0.818 0.525
20 0.101 0.241 74.62% 84.26% 80.11 83.97% 0.797 0.487
30 0.120 0.276 76.56% 84.59% 78.81% 83.03% 0.816 0.507
40 0.119 0.253 76.01% 85.40% 79.16 83.56% 0.810 0.501
50 0.106 0.254 74.85% 83.54% 77.48% 81.83% 0.825 0.506
60 0.116 0.248 75.32% 84.55% 78.44% 82.22% 0.806 0.494
Table 10: Comparison of performance on our data selection strategy with varying number of dimensions (singular values) for Subject 7. Our method has not been optimized for low-level metrics, which are shown in gray for reference. 150 images were used for all the experiments in this table.
Low-Level High-Level
Dimensions PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
10 (150) 0.112 0.272 72.60% 80.93% 74.72% 79.09% 0.840 0.509
20 (150) 0.095 0.241 69.75% 79.95% 72.92% 77.51% 0.851 0.521
30 (150) 0.102 0.258 72.06v 80.74% 74.23% 77.92% 0.843 0.526
40 (150) 0.117 0.267 73.23% 80.99% 74.20% 78.15% 0.840 0.520
50 (150) 0.105 0.220 69.18% 78.17% 70.75% 75.23% 0.860 0.533
60 (150) 0.103 0.224 69.43% 79.77% 75.07% 77.13% 0.839 0.528
Figure 15: Comparison of the performance of the image selection strategy for different subjects. (a) Inception score with varying number of dimensions (b) CLIP score with varying number of dimensions.
H.2 Impact of reference subject on performance
We conducted additional experiments (Table 11) to explore the effect of different reference subjects on reconstruction results. We use subjects 2,5 and 7 as the reference and use subject 1 as the fine-tuning subject.
Table 11: Quantitative results on fine tuning a new subject with limited data using different reference subjects. In this table we compare AAMax vs normal finetuning at 160 epochs only since AAMax significantly outperforms normal finetuning after 1 epoch of finetuning.
Low-Level High-Level
Subj1 FT From Subj2 250 Normal E160 0.007 0.268 75.55% 84.00% 75.20% 74.60% 0.862 0.520
Subj1 FT From Subj2 250 AAMAX E160 0.06 0.277 81.13% 88.99% 79.71% 78.89% 0.824 0.494
Subj1 FT From Subj5 250 Normal E160 0.094 0.239 71.54% 80.10% 73.11% 74.02% 0.875 0.537
Subj1 FT From Subj5 250 AAMAX E160 0.105 0.259 76.77% 86.20% 80.08% 80.14% 0.828 0.496
Subj1 FT From Subj7 250 Normal E160 0.010 0.266 74.96% 81.55% 71.71% 73.72% 0.874 0.528
Subj1 FT From Subj7 250 AAMAX E160 0.012 0.283 79.58% 87.78% 77.48% 77.87% 0.835 0.502
Appendix I High-Level Pipeline Training Process
@@
I.1 Training the Reference Subject
@@ I.2 Training a New Subject
Appendix J Low-Level Pipeline Training Process
@@
J.1 Training the Reference Subject
@@
$\mathcal{L}_{\text{new}}=\begin{cases}\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{fMRI, new}},\mathbf{z}_{\text{AE}}),&\text{for non-common images}\\ \mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{fMRI, new}},\mathbf{z}_{\text{AE}})+\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{adapter, new}},\mathbf{z}_{\text{adapter, ref}}),&\text{for common images}\end{cases}$
@@
Appendix K Reconstruction Results

### v2

@@ H.2 Projection to 4096 Dimensions with Stable Diffusion XL
Appendix I Common Space Empirical Analysis
@@
I.1 Eigenvector Analysis
To first investigate the structural alignment of embeddings across subjects, we analyzed the subject-wise cosine similarity of the principal eigenvectors in the embedding spaces produced by two techniques: the standard end-to-end training (Baseline) and our proposed approach (AAMax). For both techniques, we extracted the top five principal eigenvectors from the test data for each subject and computed the cosine similarity between the eigenvectors from different subjects.
@@ I.1 Eigenvector Analysis
I.2 Mean Squared Error
We compare the end-to-end training approach with AAMax by evaluating the MSE between subjects using shared images after training as shown in Figure 12. In the end-to-end model, we observe relatively high MSE between subjects, indicating poor alignment in the common space. In contrast, the MSE for AAMax is an order of magnitude lower, demonstrating a much stronger alignment of subject-specific embeddings. This result suggests that AAMax significantly enhances alignment across subjects, making it far more effective than the normal approach in bringing embeddings closer together in the common space.
Figure 12: Comparing Mean Squared Error for the shared images across subjects in both training techniques. The MSE is compared after training is completed in both techniques. MSE for AAMax is an order of magnitude lower than the baseline
I.3 k-Nearest Neighbor Consistency
We also evaluate the k-nearest neighbors (k-NN) consistency (Figure 13) between subjects for shared images after training, using 50 nearest neighbors as the metric. In the end-to-end model, the k-NN consistency is relatively low, indicating that the nearest neighbors of embeddings from one subject do not align well with those from another subject. However, with AAMax, the k-NN consistency significantly improves, demonstrating much better alignment of subject-specific embeddings. AAMax consistently outperforms the normal approach, highlighting its effectiveness in preserving semantic relationships across subjects in the common space.
Figure 13: Comparing kNN consistency @ 50 for the shared images across subjects in both training techniques. The kNN-C is compared after training is completed in both techniques. AAMax significantly outperforms the baseline.
Appendix J Emergence of Common Concepts: Additional Details
The existence of a semantic common representation space is suggested from the emergence of semantic categories in this space. For this, we extracted the most significant left singular vectors from the ridge regression weight matrix. This matrix was extracted from the pre-trained MindEye2 [Scotti et al., 2024] model which was trained using the 1024-dimensional common space. Subsequently, we projected the 1024-dimensional embeddings of each image onto these vectors and sampled from the two extremes (set 1 and set 3) and the middle (set 2). In order to obtain an objective classification, we presented these sets of images to GPT-4o [Achiam et al., 2023] and asked it to find the semantic categories that are well represented in set 1, somewhat represented in set 2 and not represented in set 3. Results for the first five singular vectors are shown in Table 11. As is evident, images in the principal dimensions of the representation space have a common theme.
@@ I.2 Mean Squared Error
Previous studies, such as those by Huth et al. [2012], have indicated that animal categories, including humans, are represented distinctly from non-animal categories within the brain. In our study, we observed that the common representational space also preserves these semantic properties. As demonstrated in Figure 14, after projecting the fMRI signals of Subject 2 into the common space and onto the first two singular vectors, we can see the animal vs. non-animal distinction, consistent with previous hypotheses. Extending this analysis across additional dimensions, we observed similar separations between other hypothesized semantic categories of "Social interaction," "man-made vs nature," and "biological vs non-biological."
@@
To determine if dominant semantic categories were present within these sets, we employed ChatGPT with the query: “Is there a dominant semantic category to which the majority of images in a set belong?” The methodology for defining semantic categories was consistent with that used in Huth et al. [2012]. Table 11 illustrates our findings.
After projecting the fMRI signals of individual images from Subject 2 into the common space and onto the first singular vectors, we observe that certain concepts are strongly represented along these dimensions defined by the singular vectors, aligning with previous hypotheses (Figures 15, 16).
@@
Appendix K NP-Hardness Proof
@@
Definition K.1 (Bin-Mapping Problem).
@@
K.2 Reduction from Set Cover
@@
Definition K.2 (Set Cover Problem).
@@
K.3 Proof of Approximation Ratio
In this section, we prove that the gap function $\text{Gap}(S)$, minimized by our greedy heuristic, satisfies the submodularity property. Submodularity guarantees the diminishing returns property, which forms the basis for the $1-\frac{1}{e}$ approximation ratio achieved by the greedy algorithm for submodular optimization  [Fisher et al., 1978].
K.3.1 Gap Function Definition
@@
K.3.2 Submodularity Property
A function $f(S)$ is submodular if it satisfies the diminishing returns property  [Schrijver et al., 2003]:
@@
K.3.5 Conclusion
Since $\text{Gap}(S)$ is both submodular and monotone, the greedy algorithm applied to minimize the gap function achieves a $1-\frac{1}{e}$ approximation ratio  [Fisher et al., 1978].
K.4 Greedy Image Selection Algorithm
@@
Appendix L Training Time
To show the efficiency of adapter alignment we computed the run time per epoch for each step of the training process. We show the AAMax takes very little extra train time to perform step 1 of the process and End-to-end fine-tuning has essentially the exact same runtime between the baseline method and AAMax. Our results are presented in Table 12.
Table 12: Comparison of Training Times for Baseline and AAMax Training Methods. Numbers are averaged over all 3 fine tune subjects (2,5,7). The Adapters are first allowed to overfit by training them upto 400 epochs using MSE loss. This takes around a minute. Then end-to-end alignment is performed. It is important to note that besides the addition of the MSE Loss at the adapter level, both pipelines are running the same architecture and data and performance is thus very similar. The final time given is the time to train up to 160 epochs. But AAMax training can outperform 160 epochs of Baseline training with just one epoch of end-to-end alignment.
@@
Baseline - 221.4 seconds 590.4 minutes 590.4 minutes
@@
Appendix M High-Level Pipeline Training Process
@@
M.1 Training the Reference Subject
@@
Appendix N Low-Level Pipeline Training Process
@@
N.1 Training the Reference Subject
@@
[…]
@@
Appendix O Reconstruction Results